# A Multi-fidelity Estimator of the Expected Information Gain for Bayesian Optimal Experimental Design

### v1

@@ Abstract
Optimal experimental design (OED) is a framework that leverages a mathematical model of the experiment to identify optimal conditions for conducting the experiment. Under a Bayesian approach, the design objective function is typically chosen to be the expected information gain (EIG). However, EIG is intractable for nonlinear models and must be estimated numerically. Estimating the EIG generally entails some variant of Monte Carlo sampling, requiring repeated data model and likelihood evaluations—each involving solving the governing equations of the experimental physics—under different sample realizations. This computation becomes impractical for high-fidelity models.
We introduce a novel multi-fidelity EIG (MF-EIG) estimator under the approximate control variate (ACV) framework. This estimator is unbiased with respect to the high-fidelity mean, and minimizes variance under a given computational budget. We achieve this by first reparameterizing the EIG so that its expectations are independent of the data models, a requirement for compatibility with ACV. We then provide specific examples under different data model forms, as well as practical enhancements of sample size optimization and sample reuse techniques. We demonstrate the MF-EIG estimator in two numerical examples: a nonlinear benchmark and a turbulent flow problem involving the calibration of shear-stress transport turbulence closure model parameters within the Reynolds-averaged Navier–Stokes model. We validate the estimator’s unbiasedness and observe one- to two-orders-of-magnitude variance reduction compared to existing single-fidelity EIG estimators.
@@ 1 Introduction
The application of ACV to OED problems remains relatively unexplored. The most closely related studies involve multilevel MC (MLMC) [19], which has also been shown to be a special case of the ACV family of estimators. In MLMC, EIG estimators are constructed where the level is controlled by the number of inner-loop samples in the NMC estimator [21] or the discretization resolution of the governing differential equations [4]. Importantly, MLMC relies on specific assumptions about the cost-accuracy relationships between models and their convergence towards a “truth model” in the limit. In contrast, ACV offers greater flexibility, as it can accommodate any ensemble of models regardless of whether their cost-accuracy relationships are known. ACV thus can incorporate diverse and flexible mixes of models, including those derived from simplified physics, different numerical methods, or data-driven approximations. In comparison, MLMC is limited to model ensembles that satisfy specific cost-accuracy and convergence assumptions, such as those based on coarsened discretizations or relaxed convergence criteria. Furthermore, ACV directly minimizes error towards the high-fidelity mean—with respect to this goal, MLMC is generally suboptimal compared to other ACV estimators.
@@ 2.1 OED problem statement
$p(\theta|y,\xi)=\frac{p(y|\theta,\xi)\,p(\theta)}{p(y|\xi)},$ (1)
@@
$U(\xi):=\mathbb{E}_{Y,\Theta|\xi}\left[u(\xi,Y,\Theta)\right],$ (2)
@@
$u(\xi,y,\theta)=\log\left[\frac{p(\theta|y,\xi)}{p(\theta)}\right]=\log\left[\frac{p(y|\theta,\xi)}{p(y|\xi)}\right],$ (3)
@@
$U(\xi)=\mathbb{E}_{Y,\Theta|\xi}\left[\log\left[\frac{p(\Theta|Y,\xi)}{p(\Theta)}\right]\right]=\int_{Y}\int_{\Theta}p(y,\theta|\xi)\,\log\left[\frac{p(\theta|y,\xi)}{p(\theta)}\right]\,\text{d}\theta\,\text{d}y.$ (4)
@@
$u^{\text{div}}(\xi,y,\theta)=D_{\text{KL}}\left(p_{\Theta|y,\xi}\,||\,p_{\Theta}\right)=\int_{\widetilde{\Theta}}p(\tilde{\theta}|y,\xi)\,\log\left[\frac{p(\tilde{\theta}|y,\xi)}{p(\tilde{\theta})}\right]\text{d}\tilde{\theta}=u^{\text{div}}(\xi,y),$ (5)
@@ 2.2 Nested Monte Carlo estimator for the EIG
$U(\xi)$ $=\int_{Y}\int_{\Theta}p(y,\theta|\xi)\,\log\left[\frac{p(y|\theta,\xi)}{p(y|\xi)}\right]\,\text{d}\theta\,\text{d}y$ (7)
$\approx\frac{1}{N_{\text{out}}}\sum^{N_{\text{out}}}_{i=1}\left\{\log\left[p(y^{(i)}|\theta^{(i)},\xi)\right]-\log\left[p(y^{(i)}|\xi)\right]\right\}$ (8)
$\approx\frac{1}{N_{\text{out}}}\sum^{N_{\text{out}}}_{i=1}\left\{\log\left[p(y^{(i)}|\theta^{(i)},\xi)\right]-\log\left[\frac{1}{N_{\text{in}}}\sum^{N_{\text{in}}}_{j=1}p(y^{(i)}|\theta^{(i,j)},\xi)\right]\right\}:=\hat{U}^{\text{NMC}}(\xi),$ (9)
where the initial equality results from transforming (4) to its likelihood-to-evidence log density ratio version, and the approximation expressions follow from standard MC estimations; $\theta^{(i)}$ are independent and identically distributed (i.i.d.) samples from the prior $p(\theta)$, and $y^{(i)}$ are corresponding independent samples from the likelihood $p(y|\theta^{(i)},\xi)$; and $N_{\text{out}}$ and $N_{\text{in}}$ are the number of MC samples in the outer and inner loops, respectively. The complexity of the NMC estimator at a given design $\xi$, in terms of likelihood operations, is $\mathcal{O}(N_{\text{out}}N_{\text{in}})$. The estimator is biased with respect to the exact EIG under finite $N_{\text{in}}$, but is asymptotically unbiased in the limit as $N_{\text{in}}\to\infty$. Detailed analysis of the NMC estimator can be found in [41, 3, 37].
@@ 3 Background on approximate control variates
$\mathbb{E}\left[Q\right]=\mathbb{E}_{Z}\left[f(Z)\right]\approx\hat{Q}(z):=\frac{1}{N}\sum^{N}_{j=1}f(z^{(j)}),$ (10)
@@
$\hat{Q}^{\text{CV}}(z;\alpha):=\hat{Q}_{0}(z)+\sum^{M}_{m=1}\alpha_{m}\left(\hat{Q}_{m}(z)-\mu_{m}\right),$ (11)
@@
$\tilde{Q}(z;\alpha,\mathcal{A}):=\hat{Q}_{0}(z_{0})+\sum^{M}_{m=1}\alpha_{m}\left(\hat{Q}_{m}(z^{\ast}_{m})-\hat{\mu}_{m}(z_{m})\right)=\hat{Q}_{0}(z_{0})+\sum^{M}_{m=1}\alpha_{m}\left(\hat{Q}_{m}(z^{\ast}_{m})-\hat{Q}_{m}(z_{m})\right),$ (12)
where now we have further allocated the input sample vector $z$ into subsets $z_{0}\neq\emptyset$, $z_{m}$, and $z^{\ast}_{m}$ such that their union is $z$, and hence the MC estimators $\hat{Q}_{m}$ all have different sample sizes corresponding to their specific input arguments. There is no strict relationship required between these sample subsets, i.e., the intersections between subsets can be empty or non-empty. Information regarding the sample allocation, including which samples are shared among different subsets and the size of each subset, is encoded in $\mathcal{A}$. This contrasts with CV, where an identical sample set is used across all $m$.
@@
$\tilde{Q}(z;\alpha,\mathcal{A})=\hat{Q}_{0}+\alpha^{\intercal}\Delta,$ (13)
@@
$\alpha^{\ast}(\mathcal{A})=-\mathbb{C}\text{ov}[\Delta,\Delta]^{-1}\mathbb{C}\text{ov}[\Delta,\hat{Q}_{0}],$ (14)
@@
$\mathbb{V}\text{ar}[\tilde{Q}(Z;\alpha^{\ast},\mathcal{A})]=\mathbb{V}\text{ar}[\hat{Q}_{0}]-\mathbb{C}\text{ov}[\Delta,\hat{Q}_{0}]^{\intercal}\mathbb{C}\text{ov}[\Delta,\Delta]^{-1}\mathbb{C}\text{ov}[\Delta,\hat{Q}_{0}].$ (15)
The sample allocation $\mathcal{A}$ determines how much computational workload is distributed to each $\hat{Q}_{m}$. In fact, both MFMC [33] and MLMC [19] are special cases of ACV that only differ in the family of possible $\mathcal{A}$ choices. For example, MFMC restricts each $z_{m}^{\ast}=z_{m-1}$ and $z_{m}\subset z_{k}$ for $m<k$. In [22], the authors generalize these rules to enable additional families of ACV estimators, reaching 16 parametric families in [8]. They further derive an expression for the variance of the ACV estimator, allowing one to compare and optimize across different allocations:
$\min_{\mathcal{A}\in\mathbb{A}}$ $\quad\mathbb{V}\text{ar}[\tilde{Q}(Z;\alpha^{\ast},\mathcal{A})]$ (16)
subject to $\quad\mathcal{W}(w,\mathcal{A})\leq w_{\text{budget}},$ (17)
@@ 4.1 Reparameterization of the EIG
The ACV framework assumes that the cost of generating input samples $z$ is negligible relative to the model evaluation costs, i.e., of $f_{m}(z)$. However, this assumption does not hold when MC is directly applied to “discretize” the expectation of the expected utility in (2), such as for the NMC estimator in (9). In these cases, sampling $Z=\{Y,\Theta\}$ involves drawing a sample of $Y$ from the likelihood, which requires evaluating the data model that simulates the experimental process. Mathematically, the data model can be written as $Y=h(\mathcal{E};\Theta,\xi)$, where $\mathcal{E}$ represents the random data noise that is assumed to be independent of $\Theta$, simple to sample, and whose probability density is easy to evaluate. A popular example is an additive independent noise structure with $Y=g(\Theta,\xi)+\mathcal{E}$, where $g$ is a deterministic forward model (e.g., governed by a system of differential equations) and $\mathcal{E}$ is a Gaussian random noise; generating a sample of $Y$ thus requires evaluating $g$ (e.g., solving the differential equations), a computationally expensive step. To satisfy the assumption for ACV, we perform a change of variables to the EIG such that the new variables of expectation can be sampled without requiring any data model.
@@
$p(y|\theta,\xi)=p_{\mathcal{E}|\xi}\left(h^{-1}(y;\theta,\xi)\big{|}\xi\right)\,\big{\lvert}J^{-1}(y;\theta,\xi)\big{\rvert}=p(\epsilon|\xi)\,\big{\lvert}J^{-1}(h(\epsilon;\theta,\xi);\theta,\xi)\big{\rvert},$ (18)
@@
$p(y|\xi)=\int_{\widetilde{\Theta}}p(y|\tilde{\theta},\xi)\,p(\tilde{\theta})\,\text{d}{\tilde{\theta}}=\int_{\widetilde{\Theta}}p_{\mathcal{E}|\xi}(\tilde{\epsilon}|\xi)\,\big{\lvert}J^{-1}(y;\tilde{\theta},\xi)\big{\rvert}\,p(\tilde{\theta})\,\text{d}\tilde{\theta},$ (19)
@@
$=\int_{\mathcal{E}}\int_{\Theta}p(\theta)\,p(\epsilon|\xi)\,\big{\lvert}J^{-1}(h(\epsilon;\theta,\xi);\theta,\xi)\big{\rvert}\,\log\left[\frac{p(\epsilon|\xi)\,\big{\lvert}J^{-1}(h(\epsilon;\theta,\xi);\theta,\xi)\big{\rvert}}{\int_{\widetilde{\Theta}}p_{\mathcal{E}|\xi}(\tilde{\epsilon}|\xi)\,\big{\lvert}J^{-1}(h({\epsilon};{\theta},\xi);\tilde{\theta},\xi)\big{\rvert}\,p(\tilde{\theta})\,\text{d}\tilde{\theta}}\right]\,\text{d}\theta\,\big{\lvert}J(\epsilon;\theta,\xi)\big{\rvert}\,\text{d}\epsilon$
$=\int_{\mathcal{E}}\int_{\Theta}p(\epsilon,\theta|\xi)\,\log\left[\frac{p(\epsilon|\xi)\,\big{\lvert}J^{-1}(h(\epsilon;\theta,\xi);\theta,\xi)\big{\rvert}}{\int_{\widetilde{\Theta}}p_{\mathcal{E}|\xi}(\tilde{\epsilon}|\xi)\,\big{\lvert}J^{-1}(h(\epsilon;\theta,\xi);\tilde{\theta},\xi)\big{\rvert}\,p(\tilde{\theta})\,\text{d}\tilde{\theta}}\right]\,\text{d}\theta\,\text{d}\epsilon$ (20)
$=\mathbb{E}_{\mathcal{E},\Theta|\,\xi}\left[\log\left[\frac{p(\mathcal{E}|\xi)\,\big{\lvert}J^{-1}(h(\mathcal{E};\Theta,\xi);\Theta,\xi)\big{\rvert}}{\int_{\widetilde{\Theta}}p_{\mathcal{E}|\xi}(\tilde{\mathcal{E}}|\xi)\,\big{\lvert}J^{-1}(h({\mathcal{E}};{\Theta},\xi);\tilde{\theta},\xi)\big{\rvert}\,p(\tilde{\theta})\,\text{d}\tilde{\theta}}\right]\right]=\mathbb{E}_{\mathcal{E},\Theta|\,\xi}\left[u_{\mathcal{E}}(\xi,\mathcal{E},\Theta)\right],$ (21)
@@
$u_{\mathcal{E}}(\xi,\epsilon,\theta):=\log\left[\frac{p(\epsilon|\xi)\,\big{\lvert}J^{-1}(h(\epsilon;\theta,\xi);\theta,\xi)\big{\rvert}}{\int_{\widetilde{\Theta}}p_{\mathcal{E}|\xi}(\tilde{\epsilon}|\xi)\,\big{\lvert}J^{-1}(h({\epsilon};{\theta},\xi);\tilde{\theta},\xi)\big{\rvert}\,p(\tilde{\theta})\,\text{d}\tilde{\theta}}\right]$ (22)

### v2

@@ Abstract
Optimal experimental design (OED) is a framework that leverages a mathematical model of the experiment to identify optimal conditions for conducting the experiment. Under a Bayesian approach, the design objective function is typically chosen to be the expected information gain (EIG). However, EIG is intractable for nonlinear models and must be estimated numerically. Estimating the EIG generally entails some variant of Monte Carlo sampling, requiring repeated data model and likelihood evaluations—each involving solving the governing equations of the experimental physics—under different sample realizations. This computation becomes impractical for high-fidelity models. We introduce a novel multi-fidelity EIG (MF-EIG) estimator under the approximate control variate (ACV) framework. This estimator is unbiased with respect to the high-fidelity mean, and minimizes variance under a given computational budget. We achieve this by first reparameterizing the EIG so that its expectations are independent of the data models, a requirement for compatibility with ACV. We then provide specific examples under different data model forms, as well as practical enhancements of sample size optimization and sample reuse techniques. We demonstrate the MF-EIG estimator in two numerical examples: a nonlinear benchmark and a turbulent flow problem involving the calibration of shear-stress transport turbulence closure model parameters within the Reynolds-averaged Navier–Stokes model. We validate the estimator’s unbiasedness and observe one- to two-orders-of-magnitude variance reduction compared to existing single-fidelity EIG estimators.
keywords:
Uncertainty quantification, Approximate control variates, Variance reduction, Turbulence closure
{MSCcodes}
62K05, 62C10, 62C05
@@ 1 Introduction
The application of ACV to OED problems remains relatively unexplored. The most closely related studies involve multilevel MC (MLMC) [19], which has also been shown to be a special case of the ACV family of estimators. In MLMC, EIG estimators are constructed where the level is controlled by the number of inner-loop samples in the NMC estimator [21] or the discretization resolution of the governing differential equations [4]. Importantly, MLMC relies on specific assumptions about the cost-accuracy relationships between models and theirconvergence towards a “truth model” in the limit. In contrast, ACV offers greater flexibility, as it can accommodate any ensemble of models regardless of whether their cost-accuracy relationships are known. ACV thus can incorporate diverse and flexible mixes of models, including those derived from simplified physics, different numerical methods, or data-driven approximations. In comparison, the theory surrounding MLMC is limited to model ensembles that satisfy specific cost-accuracy and convergence assumptions, such as those based on coarsened discretizations or relaxed convergence criteria. Furthermore, ACV directly minimizes error towards the high-fidelity mean—with respect to this goal, MLMC is generally suboptimal compared to other ACV estimators.
@@ 2.1 OED problem statement
(1) $p(\theta|y,\xi)=\frac{p(y|\theta,\xi)\,p(\theta)}{p(y|\xi)},$
@@
(2) $U(\xi):=\mathbb{E}_{Y,\Theta|\xi}\left[u(\xi,Y,\Theta)\right],$
@@
(3) $u(\xi,y,\theta)=\log\left[\frac{p(\theta|y,\xi)}{p(\theta)}\right]=\log\left[\frac{p(y|\theta,\xi)}{p(y|\xi)}\right],$
@@
(4) $U(\xi)=\mathbb{E}_{Y,\Theta|\xi}\left[\log\left[\frac{p(\Theta|Y,\xi)}{p(\Theta)}\right]\right]=\int_{Y}\int_{\Theta}p(y,\theta|\xi)\,\log\left[\frac{p(\theta|y,\xi)}{p(\theta)}\right]\,\text{d}\theta\,\text{d}y.$
@@
(5) $u^{\text{div}}(\xi,y,\theta)=D_{\text{KL}}\left(p_{\Theta|y,\xi}\,||\,p_{\Theta}\right)=\int_{\widetilde{\Theta}}p(\tilde{\theta}|y,\xi)\,\log\left[\frac{p(\tilde{\theta}|y,\xi)}{p(\tilde{\theta})}\right]\text{d}\tilde{\theta}=u^{\text{div}}(\xi,y),$
@@ 2.2 Nested Monte Carlo estimator for the EIG
(7) $U(\xi)$ $=\int_{Y}\int_{\Theta}p(y,\theta|\xi)\,\log\left[\frac{p(y|\theta,\xi)}{p(y|\xi)}\right]\,\text{d}\theta\,\text{d}y$
(8) $\approx\frac{1}{N_{\text{out}}}\sum^{N_{\text{out}}}_{i=1}\left\{\log\left[p(y^{(i)}|\theta^{(i)},\xi)\right]-\log\left[p(y^{(i)}|\xi)\right]\right\}$
(9) $\approx\frac{1}{N_{\text{out}}}\sum^{N_{\text{out}}}_{i=1}\left\{\log\left[p(y^{(i)}|\theta^{(i)},\xi)\right]-\log\left[\frac{1}{N_{\text{in}}}\sum^{N_{\text{in}}}_{j=1}p(y^{(i)}|\theta^{(i,j)},\xi)\right]\right\}:=\hat{U}^{\text{NMC}}(\xi),$
where the initial equality results from transforming (4) to its likelihood-to-evidence log density ratio version, and the approximation expressions follow from standard MC estimations; $\theta^{(i)}$ are independent and identically distributed (i.i.d.) samples from the prior $p(\theta)$, and $y^{(i)}$ are corresponding independent samples from the likelihood $p(y|\theta^{(i)},\xi)$; and $N_{\text{out}}$ and $N_{\text{in}}$ are the number of MC samples in the outer and inner loops, respectively. The complexity of the NMC estimator at a given design $\xi$, in terms of likelihood operations, is $\mathcal{O}(N_{\text{out}}N_{\text{in}})$. The estimator is biased with respect to the exact EIG under finite $N_{\text{in}}$, but is asymptotically unbiased in the limit as $N_{\text{in}}\to\infty$. Detailed analysis of the NMC estimator can be found in [41, 3, 37]. We note that alternative estimators—such as those referenced in Section 1—can also be employed; however, they may introduce additional bias, for example when relying on Gaussian approximations or surrogate models.
@@ 3 Background on approximate control variates
(10) $\mathbb{E}\left[Q\right]=\mathbb{E}_{Z}\left[f(Z)\right]\approx\hat{Q}(z):=\frac{1}{N}\sum^{N}_{j=1}f(z^{(j)}),$
@@
(11) $\hat{Q}^{\text{CV}}(z;\alpha):=\hat{Q}_{0}(z)+\sum^{M}_{m=1}\alpha_{m}\left(\hat{Q}_{m}(z)-\mu_{m}\right),$
@@
(12) $\tilde{Q}(z;\alpha,\mathcal{A}):=$ $\,\,\hat{Q}_{0}(z_{0})+\sum^{M}_{m=1}\alpha_{m}\left(\hat{Q}_{m}(z^{\ast}_{m})-\hat{\mu}_{m}(z_{m})\right)$
$=$ $\,\,\hat{Q}_{0}(z_{0})+\sum^{M}_{m=1}\alpha_{m}\left(\hat{Q}_{m}(z^{\ast}_{m})-\hat{Q}_{m}(z_{m})\right),$
where now we have further allocated the input sample vector $z$ into subsets $z_{0}\neq\emptyset$, $z_{m}$, and $z^{\ast}_{m}$ such that their union is $z$, and hence the MC estimators $\hat{Q}_{m}$ all have different sample sizes corresponding to their specific input arguments. There is no strict relationship required between these sample subsets, i.e., the intersections between subsets can be empty or non-empty. Information regarding the sample allocation, including which samples are shared among different subsets and the size of each subset, is encoded in $\mathcal{A}$. This contrasts with CV, where an identical sample set is used across all $m$. Most tools also support the exclusion of specific models by setting their corresponding sample allocations to $\emptyset$, as is implemented in the ‘Automatic Model Selection’ feature of MXMCPy [9]. This enables users to include all available models without a priori selection, with the only tradeoff being a slight increase in computation time when solving the optimal sample allocation.
@@
(13) $\tilde{Q}(z;\alpha,\mathcal{A})=\hat{Q}_{0}+\alpha^{\intercal}\Delta,$
@@
(14) $\alpha^{\ast}(\mathcal{A})=-\mathbb{C}\text{ov}[\Delta,\Delta]^{-1}\mathbb{C}\text{ov}[\Delta,\hat{Q}_{0}],$
@@
(15) $\mathbb{V}\text{ar}[\tilde{Q}(Z;\alpha^{\ast},\mathcal{A})]=\mathbb{V}\text{ar}[\hat{Q}_{0}]-\mathbb{C}\text{ov}[\Delta,\hat{Q}_{0}]^{\intercal}\mathbb{C}\text{ov}[\Delta,\Delta]^{-1}\mathbb{C}\text{ov}[\Delta,\hat{Q}_{0}].$
The sample allocation $\mathcal{A}$ determines how much computational workload is distributed to each $\hat{Q}_{m}$. In fact, both MFMC [33] and MLMC [19] are special cases of ACV that differ in the family of possible $\mathcal{A}$ choices and how $\alpha$ is set. For example, MFMC restricts each $z_{m}^{\ast}=z_{m-1}$ and $z_{m}\subset z_{k}$ for $m<k$, and MLMC uses $\alpha_{m}=-1$ for all $m=1,\ldots,M$. We also note that while the MLMC estimator is in fact an ACV estimator by structure, the theory [19] regarding its error convergence and computational complexity bounds only apply under multilevel assumptions on the model ensemble used. In [22], the authors generalize these rules to enable additional families of ACV estimators, reaching 16 parametric families in [8]. They further derive an expression for the variance of the ACV estimator, allowing one to compare and optimize across different allocations:
(16) $\min_{\mathcal{A}\in\mathbb{A}}$ $\quad\mathbb{V}\text{ar}[\tilde{Q}(Z;\alpha^{\ast},\mathcal{A})]$
(17) subject to $\quad\mathcal{W}(w,\mathcal{A})\leq w_{\text{budget}},$
@@ 4.1 Reparameterization of the EIG
The ACV framework assumes that the cost of generating input samples $z$ is negligible relative to the model evaluation costs, i.e., of $f_{m}(z)$. However, this assumption does not hold when MC is directly applied to “discretize” the expectation of the expected utility in (2), such as for the NMC estimator in (9). In these cases, sampling $Z=\{Y,\Theta\}$ involves drawing a sample of $Y$ from the likelihood, which requires evaluating the data model that simulates the experimental process. Mathematically, the data model can be written as $Y=h(\mathcal{E};\Theta,\xi)$, where $\mathcal{E}$ represents the random data noise that is assumed to be independent of $\Theta$, simple to sample, and whose probability density is easy to evaluate. A common example of $h$ is an additive independent noise model in the form $Y=g(\Theta,\xi)+\mathcal{E}$, where $g$ is a deterministic forward model (e.g., governed by a system of differential equations) and $\mathcal{E}$ is a Gaussian random noise. However, we emphasize that the function $h$ in our formulation is general and not restricted to the additive noise case. To satisfy the assumption for ACV, we perform a change of variables to the EIG such that the new variables of expectation can be sampled without requiring any data model.
@@
(18) $p(y|\theta,\xi)=p_{\mathcal{E}|\xi}\left(h^{-1}(y;\theta,\xi)\big{|}\xi\right)\,\big{\lvert}J^{-1}(y;\theta,\xi)\big{\rvert}=p(\epsilon|\xi)\,\big{\lvert}J^{-1}(h(\epsilon;\theta,\xi);\theta,\xi)\big{\rvert}$,
@@
(19) $p(y|\xi)=\int_{\widetilde{\Theta}}p(y|\tilde{\theta},\xi)\,p(\tilde{\theta})\,\text{d}{\tilde{\theta}}=\int_{\widetilde{\Theta}}p_{\mathcal{E}|\xi}(\tilde{\epsilon}|\xi)\,\big{\lvert}J^{-1}(y;\tilde{\theta},\xi)\big{\rvert}\,p(\tilde{\theta})\,\text{d}\tilde{\theta}$,
@@
$=\int_{\mathcal{E}}\int_{\Theta}p(\theta)\,p(\epsilon|\xi)\,\big{\lvert}J^{-1}(h(\epsilon;\theta,\xi);\theta,\xi)\big{\rvert}\,$
$\hskip 65.00009pt\times\log\left[\frac{p(\epsilon|\xi)\,\big{\lvert}J^{-1}(h(\epsilon;\theta,\xi);\theta,\xi)\big{\rvert}}{\int_{\widetilde{\Theta}}p_{\mathcal{E}|\xi}(\tilde{\epsilon}|\xi)\,\big{\lvert}J^{-1}(h({\epsilon};{\theta},\xi);\tilde{\theta},\xi)\big{\rvert}\,p(\tilde{\theta})\,\text{d}\tilde{\theta}}\right]\,\text{d}\theta\,\big{\lvert}J(\epsilon;\theta,\xi)\big{\rvert}\,\text{d}\epsilon$
(20) $=\int_{\mathcal{E}}\int_{\Theta}p(\epsilon,\theta|\xi)\,\log\left[\frac{p(\epsilon|\xi)\,\big{\lvert}J^{-1}(h(\epsilon;\theta,\xi);\theta,\xi)\big{\rvert}}{\int_{\widetilde{\Theta}}p_{\mathcal{E}|\xi}(\tilde{\epsilon}|\xi)\,\big{\lvert}J^{-1}(h(\epsilon;\theta,\xi);\tilde{\theta},\xi)\big{\rvert}\,p(\tilde{\theta})\,\text{d}\tilde{\theta}}\right]\,\text{d}\theta\,\text{d}\epsilon$
(21) $=\mathbb{E}_{\mathcal{E},\Theta|\,\xi}\left[\log\left[\frac{p(\mathcal{E}|\xi)\,\big{\lvert}J^{-1}(h(\mathcal{E};\Theta,\xi);\Theta,\xi)\big{\rvert}}{\int_{\widetilde{\Theta}}p_{\mathcal{E}|\xi}(\tilde{\mathcal{E}}|\xi)\,\big{\lvert}J^{-1}(h({\mathcal{E}};{\Theta},\xi);\tilde{\theta},\xi)\big{\rvert}\,p(\tilde{\theta})\,\text{d}\tilde{\theta}}\right]\right]=\mathbb{E}_{\mathcal{E},\Theta|\,\xi}\left[u_{\mathcal{E}}(\xi,\mathcal{E},\Theta)\right],$
@@
(22) $u_{\mathcal{E}}(\xi,\epsilon,\theta):=\log\left[\frac{p(\epsilon|\xi)\,\big{\lvert}J^{-1}(h(\epsilon;\theta,\xi);\theta,\xi)\big{\rvert}}{\int_{\widetilde{\Theta}}p_{\mathcal{E}|\xi}(\tilde{\epsilon}|\xi)\,\big{\lvert}J^{-1}(h({\epsilon};{\theta},\xi);\tilde{\theta},\xi)\big{\rvert}\,p(\tilde{\theta})\,\text{d}\tilde{\theta}}\right$]